I borrowed 1000$ at the end of the 3rd month of 2021. If the interest rate is 10% every 6 months and I have to pay monthly for 3 years, how much will I need to pay if my 1st payment is due on the 2nd year?


1
Expert's answer
2021-03-16T12:01:33-0400
"\frac {1000\times (1+0.1)^6}{36}=49.2"

Per one month if will bepayment every month.


If the first payment will be due on the 2nd year:


"\frac {49.2\times 12}{24}+49.2=73.8"